### v1

# LLM-based Discriminative Reasoning for Knowledge Graph Question Answering

Abstract
Large language models (LLMs) based on generative pre-trained Transformer have achieved remarkable performance on knowledge graph question-answering (KGQA) tasks. However, LLMs often produce ungrounded subgraph planning or reasoning results in KGQA due to the hallucinatory behavior brought by the generative paradigm, which may hinder the advancement of the LLM-based KGQA model. To deal with the issue, we propose a novel LLM-based Discriminative Reasoning (LDR) method to explicitly model the subgraph retrieval and answer inference process. By adopting discriminative strategies, the proposed LDR method not only enhances the capability of LLMs to retrieve question-related subgraphs but also alleviates the issue of ungrounded reasoning brought by the generative paradigm of LLMs. Experimental results show that the proposed approach outperforms multiple strong comparison methods, along with achieving state-of-the-art performance on two widely used WebQSP and CWQ benchmarks  [^1] Our code and data will be released..
[…]
LLM-based Discriminative Reasoning for Knowledge Graph Question Answering
Mufan Xu, Kehai Chen, Xuefeng Bai, Muyun Yang, Tiejun Zhao, Min Zhang School of Computer Science and Technology, Harbin Institute of Technology, China xmuffins0610@gmail.com, {chenkehai,baixuefeng,yangmuyun,tjzhao,zhangmin2021}@hit.edu.cn
1 Introduction
Large language models (LLMs) have shown remarkable reasoning capabilities in KGQA task Yu et al. (2022); Huang and Chang (2023); Wang et al. (2023b), especially the feasibility to prompt the LLMs to generate searching and reasoning results through the LLMs’ built-in knowledge. Typically, based on the given question, LLMs can be prompted to provide a plan for the question-related subgraph through one-time generation. After retrieving the subgraph, LLMs can directly generate the answers along with the reasoning steps using the subgraph as context. Utilizing internal knowledge or reasoning ability distilled from stronger models like GPT-4, the generative KGQA model can effectively conduct knowledge graph reasoning, along with achieving state-of-the-art performance on the KGQA tasks Mondorf and Plank (2024); LUO et al. (2024); Sun et al. (2024a).
Figure 1: Example of previous generation-based methods and discriminative method proposed in this paper. The generation-based methods tend to generate unsupported or redundant subgraphs and reasoning results, while discriminative method address this issue by focusing only on the KG and the subgraph.
Despite their success, the generative LLMs often produce ungrounded subgraph planning or reasoning results due to the hallucinatory behavior brought by the generation process Zhang et al. (2023); Sun et al. (2024b); Pan et al. (2024), which is opposite to the deterministic characteristic of knowledge reasoning process Garcez et al. (2015); Xiong et al. (2024). As shown in Figure 1, when searching question-related subgraphs, generation-based methods come with not existed path “1->5->3” due to hallucinatory planning (upper left), or retrieve redundant paths at one step “1->2; 1->3; 1->4” as a compensate to the generation uncertainty (upper middle). These issues lead to a noisy retrieved subgraph, which disperses the LLM’s attention during the reasoning process. When conducting answer inference on the retrieved subgraph, the generation-based methods may generate unreasonable reasoning steps “since r3" as inference chain or even entity “4” that do not exist in the subgraphs as answers (bottom left). As a result, the hallucinatory behavior of the generative LLMs hinders the performance of KGQA.
To address this issue, we propose a novel LLM-based Discriminative Reasoning (LDR) method to explicitly model the subgraph retrieval and answer inference process in the KGQA task. To this end, we reformulate the KGQA process into a discriminative reasoning framework, where KGQA is reformulated as three subtasks including subgraph searching, subgraph pruning, and answer inference. Thus, an LLM-based discriminative reasoning network is designed to model subgraph retrieval, pruning, and answer inference, thereby better achieving the KGQA. In summary, our main contributions are as follows:
  • We propose a discriminative reasoning framework to model the reasoning skills in KGQA. To the best of our knowledge, this is the first work to model different reasoning skills explicitly for enhancing the KGQA.
  • The proposed LDR organically combines the existing generative LLMs and our discriminative reasoning framework to advance the progress of KGQA tasks.
  • Experimental results demonstrate that LDR achieved state-of-the-art performance on two widely used benchmarks.
2 Discriminative Reasoning Framework
Figure 2: The proposed LDR of KGQA. Start from the question (bottom left) with a given starting entity "Coronation Street", LDR sequentially conducts subgraph retrieval, subgraph pruning, and answer inference. Then LDR automatically uses the reasoning results to prune $G_{k}$ and then retrieve the answers from it. In this figure, the node’s color in subgraph $G_{k}$ (middle center) represents its position in subgraph structure $S_{k}$ (top center).
In this section, we propose a novel discriminative reasoning framework for KGQA which reformulates the KGQA task into three subtasks, including question-related subgraph searching, question-related subgraph pruning, and answer inference. Here we first formulate the KGQA task.
Formulation of KGQA task. Given a question $Q$ and the knowledge graph entities $E$ contained in the question, the KGQA task asks the model to recall golden answers $A_{gold}$ as much as possible. The model has to retrieve question-related subgraphs from the knowledge graph $KG$ and infer the right answer based on the subgraphs.
The knowledge graph $KG$ used in this paper is Freebase  [^2] The two benchmarks used in this work are constructed using Freebase Bollacker et al. (2008)., which consists of knowledge triplets represented as $t=(s,p,o)$ including the subject entity $s$, the object entity $o$, and the predicate $p$ that connects these two entities. Next, we will introduce the three subtasks in detail.
2.1 Question-related subgraph Searching
Formulation Given the knowledge graph, the question-related subgraph searching subtask aims to retrieve the question-related subgraph $G_{k}$ and thereby summarizes an abstract structure $S_{k}$ for each input question.
Specifically, $G_{k}$ is the subgraph comprising only the necessary knowledge to correctly answer the question $Q$, which can be represented by a set of triplets:
$G_{k}=\{t_{i}|t_{i}=(s_{i},p_{i},o_{i})\}.$ (1)
The summarized abstract structure $S_{k}$ groups all the entities into “semantic nodes” based on their position in $G_{k}$:
$S_{k}=\{t_{i}|t_{i}=(s_{abs},p_{j},o_{abs})\},$ (2)
$s_{abs},o_{abs}\in Group(G_{k}).$
For example, if $G_{k}$ includes two triples, ($a$, $friend\_of$, $b$) and ($a$, $friend\_of$, $c$), its abstract structure $S_{k}$ is $\{entity_{1},friend\_of,entity_{2}\}.$ Here both entities $b$ and $c$ are grouped into the abstract node $entity\_2$ since they connect to the same entity $a$ with the same relation $friend\_of$. Note that $S_{k}$ only groups the entities and keeps the name of the relations.
2.2 Question-related Subgraph Pruning
Formulation Based on the abstract structure $S_{k}$, the question-related subgraph pruning task aims to map all the question-related constraint entities $C$ to nodes in $S_{k}$.
$C$ denotes the intersection of question mentioned entities $E_{question}$ and all entities in Freebase $E_{freebase}$:
$C=E_{question}\cap E_{freebase}.$ (3)
Let Node($S_{k}$) represents the set of nodes in $S_{k}$, the mapping results between entities in $C$ and nodes in $S_{k}$ is represented as:
$\{(C_{i},N_{i})|C_{i}\in C,N_{i}\in Node(S_{k})\}.$ (4)
For example, when answering the question: "What is the name of the team who won the Super Bowl in 2011?", there are two constraint entities $C_{1}$: "Super Bowl" and $C_{2}$: "2011". Now that the retrieved subgraph $G_{k}$ is always a tree rooted from the starting entity, any branches contain information against the the information in $C_{1}$ should be pruned from its root. In order to better focus on the LLM-based discriminative reasoning process, we assume that the entities mentioned in the questions have already been linked to Freebase entities through rule-based recognition methods.
2.3 Answer Inference
Formulation Given the question-related subgraph structure $S_{k}$, the answer inference sub-task aims to locate the position of the answer $A_{pos}$ in structure $S_{k}$.
The position of answers $A_{pos}$ corresponds to the position of $A_{gold}$ in $S_{k}$:
$A_{pos}=Grouped(A_{gold}),$ (5)
$A_{gold}\in Node(G_{k}),$
$A_{pos}\in Node(S_{k}).$
Once the position of the answer $A_{pos}$ is selected, the corresponding group of entities in $G_{k}$ will be regarded as the final answers.
3 LLM-based Discriminative Reasoning
Targeting three subtasks in the discriminative reasoning framework, we propose LLM-based discriminative reasoning (LDR) to explicitly complete the subtasks as illustrated in Figure 2.
3.1 Discriminative Searching
Previous methods rely on agent toolboxes add to the interaction overhead between the model and the knowledge graph and suffer from SPARQL syntax errors Cao et al. (2022); Gu et al. (2023); Jiang et al. (2024). Therefore, LDR opts for a discriminative searching approach which only observes and updates the subgraph structure $S_{k}$ during the process.
In each iteration, based on the retrieved subgraph structure $S_{k}$, the LLM selects one option from the option pool (as shown in Figure 2). Each option includes a starting node $s_{next}$ and a neighboring relation $p_{next}$, forming the next triple to search for:
$t_{next}=(s_{next},p_{next},o_{next}),$ (6)
$s_{next}\in Node(S_{k}).$
Since each node in the subgraph structure $S_{k}$ corresponds to a set of entities in $G_{k}$, conducting a search from a certain point in $S_{k}$ represents performing a search from each corresponding entity in $G_{k}$. A new node $o_{next}$ is added to $S_{k}$ along with its corresponding entities in $G_{k}$ retrieved from the knowledge graph. LDR maintains an option pool that includes all feasible triples for search. We can formulate each step of the discriminative searching strategy in LDR as:
$t_{next}=argmax(\mathbb{P}(t|S_{k},Q),t\in pool),$ (7)
where $\mathbb{P}$ represents the probability distribution over options provided by the LLM using a constrained beam search algorithm based on output logits. After each decision made based on the option pool, LDR removes the selected option from the pool and expands the available neighboring directions introduced by the new entities. An additional option, ’None’, is always available to terminate the search process.
Retention of Node Information. LDR further enriches the information contained within the subgraph structure $S_{k}$ by labeling the semantic nodes with entity types. Entities in Freebase can be classified into one of the following types: entity, topic, date, and num (details are provided in Appendix A). LDR will recognize the type of retrieved entities, and use these to name newly added nodes in $S_{k}$.
3.2 Discriminative Pruning
After obtaining the question-related subgraph $G_{k}$ along with its structure $S_{k}$, LDR maps all constraints mentioned in the question onto $S_{k}$ to perform subgraph pruning. LDR determines the positions of constraints observing only the retrieved subgraph structure $S_{k}$, thereby isolating the influence of specific entity names in the task.
To apply a constraint $C_{n}$ to the subgraph, we need three elements: the constraint entity $c_{tar}$ mentioned in the questions, the target position $c_{pos}$ for applying the constraint, and the operator $c_{opt}$ to define the type of logical resolution used for applying the constraint: $C_{n}$ = ($c_{pos}$, $c_{opt}$, $c_{tar}$). We restrict the operator $c_{opt}$ to one of the seven types $\{=,<,\leq,>,\geq,\text{min},\text{max}\}$ and combine each constraint with all possible operators and positions to form the option pool.
Based on the question $Q$, structure $S_{k}$, LDR asks the LLM to iteratively select the constraints until the LLM selects ’None’ or there is no options left:
$C_{n}=argmax(\mathbb{P}(C_{n}|C_{1}...,C_{n-1},S_{k},Q)).$ (8)
Once the pruning process is terminated by the LLM, LDR applies the constraint operations $C_{1},...,C_{n}$ to prune the subgraph $G_{k}$. The pruning process is conducted at the level of subtrees rooted from the starting node in $G_{k}$ (as shown in Figure 2), retaining only the subtrees that meet the constraints.
Method WebQSP CWQ
Hits@1 Recall F1 Hits@1 Recall F1
Llama2-7b zero-shot (Touvron et al., 2023)* 0.403 - 0.293 0.297 - 0.272
Llama3-8b zero-shot (Dubey et al., 2024)* 0.303 - 0.257 0.305 - 0.278
Qwen2.5-7b zero-shot (Yang et al., 2024)* 0.284 - 0.237 0.259 - 0.241
KV-Mem (Miller et al., 2016) 0.467 - 0.345 0.184 - 0.157
GraftNet (Sun et al., 2018) 0.664 - 0.604 0.368 - 0.327
QGG (Lan and Jiang, 2020) 0.730 - 0.738 0.369 - 0.374
NSM (He et al., 2021) 0.687 - 0.628 0.476 - 0.424
SR+NSM (Zhang et al., 2022) 0.689 - 0.641 0.502 - 0.471
SR+NSM+E2E (Zhang et al., 2022) 0.695 - 0.641 0.493 - 0.463
DECAF (DPR+FiD-3B) (Yu et al., 2022) 0.821 - 0.788 - - -
UniKGQA (Jiang et al., 2022) 0.751 - 0.702 0.507 - 0.480
KD-CoT (Wang et al., 2023a) 0.686 - 0.525 0.557 - -
ToG w/ChatGPT (Sun et al., 2024a) 0.758 - - 0.589 - -
ToG w/GPT-4 (Sun et al., 2024a) 0.826 - - 0.676 - -
KG-Agent (Jiang et al., 2024) 0.833 - 0.810 0.722 - 0.692
RoG (Top-3 relation path) (LUO et al., 2024)* 0.795 0.756 0.701 0.567 0.573 0.547
LDR (Ours) 0.840 0.860 0.845 0.802 0.837 0.820
Table 1: The results of our method compared with previous approaches on WebQSP and CWQ. Asterisk * denotes the results we reproduced. Note that the Hits@1 result reported in the original RoG paper (WebQSP 0.857, CWQ 0.626) is not calculated in the right way, see the author’s response here.
3.3 Discriminative Answer inferring
When answering questions with a large number of answers, previous generative methods often fail to capture all the correct answers, even if the reasoning steps are successfully generated. To address this problem, LDR focuses on locating the positions of answers within the subgraph structure $S_{k}$ to simultaneously retrieve all possible answers. Based on $S_{k}$, LDR determines the answer position $A_{pos}$ using:
$A_{pos}=argmax(\mathbb{P}(n|S_{k},C,Q),n\in S_{k}).$ (9)
Here $\mathbb{P}$ is also given by the LLM based on $S_{k}$,$C$ and $Q$. Positions for applying constraints can not be chosen again. All entities in $G_{k}$ corresponding to the selected position $A_{pos}$ will be listed as the answer.
3.4 Constructing Training Data
To train the model, we utilize the SPARQL queries provided in the benchmarks. Here we briefly introduce the training data used for finetuning the LLMs in LDR.
Subgraph Searching. We make use of the SPARQL data available in existing benchmarks to form the training data. In WebQSP and CWQ, each question is associated with a SPARQL query. The direct execution of this query yields the answer to the open question (for details with SPARQL decomposition, see Appendix B). We obtain the correct subgraph structure required to solve each problem by decomposing the SPARQL statements. Unlike ROG (LUO et al., 2024), in finetuning process LDR always presents the model with the correct knowledge subgraph structure rather than the shortest path starts from the question entity and ends at the answer entities.
Subgraph pruning and answer inference. To finetune the LLMs to be capable of constraint determination and answer inference, we also construct constraint/answer locating samples from the SPARQL queries in WebQSP and CWQ. The input is a complete subgraph structure with all feasible options of constraints or answer positions, the golden output is the correct position of the constraint and the answer.
Data Statistics. Using the method described above, we construct our training set with 121,023 subgraph searching samples and 46,885 subgraph pruning and answer inference samples.
4 Experiments
4.1 Datasets and Settings
Benchmarks. To evaluate the knowledge graph question-answering capability of the proposed method, we choose two widely used benchmarks, WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018) (for more details, see Appendix B).
Metrics. We choose commonly used metrics Hits@1 and F1 for the evaluation process following previous works LUO et al. (2024); Sun et al. (2024a); Jiang et al. (2024). For detailed definitions of metrics, please refer to Appendix C.
Baselines. We compare LDR with previous SOTA generation-based KGQA strategies including RoG with LLM planning and chain-of-thought reasoning strategy (LUO et al., 2024), ToG with interactive reasoning strategy (Sun et al., 2024a) and KGAgent with tool-based agent reasoning strategy (Jiang et al., 2024). We also list previous representative methods and zero-shot performances of three widely used LLMs for comparison.
Base Model. To ensure fairness in comparison, we choose Llama2-7b (Touvron et al., 2023) as the base model of LDR following RoG (LUO et al., 2024) and KGAgent (Jiang et al., 2024).
4.2 Main results
The performance of LDR on WebQSP and CWQ is presented in Table 1. The LDR method achieves state-of-the-art performance on these two benchmarks, with improvements in both Hits@1 and F1, indicating an enhanced capability of the LLM to handle KGQA tasks. Despite foregoing the possibility for the LLM to use its own knowledge to directly answer questions correctly, LDR still outperforms previous generative methods, which demonstrates the effectiveness of the proposed strategies (refer to Appendix D for details).
4.3 Enhancement of Subgraph Quality
Leaf Number 2 3 4 5 Avg.
Total Hop 1 2 3 4 5 3 4 5 4 5 6 5 6 7
Relation Recall $R_{rel}$
RoG 0.853 0.644 0.381 0.280 0.254 0.429 0.266 0.270 0.286 0.179 0.169 0.186 0.266 0.283 0.339
LDR 0.887 0.887 0.903 0.897 0.972 0.859 0.853 0.899 0.748 0.656 0.826 0.867 0.837 0.863 0.853
Minimum Graph Edit Distance $D(G_{k},G_{gold})$
RoG 0.479 2.494 3.929 5.462 7.727 3.071 3.746 5.394 1.760 4.780 5.681 5.441 8.100 10.438 4.893
LDR 0.097 0.209 0.315 0.625 0.181 0.338 1.069 1.490 1.521 3.658 3.000 1.235 1.550 1.578 1.204
Table 2: We use relation recall and minimum graph edit distance as the metrics to measure the quality of retrieved subgraphs with different type of structures. Our method consistently achieves higher relation recall and lower average edit distance than RoG using generation-based searching strategy.
To validate LDR’s enhancement on the capability of LLM to retrieve question-related subgraphs, we design two metrics, relation recall and minimum graph edit distance, to measure the difference between the retrieved subgraph $G_{k}$ and the golden subgraph $G_{gold}$ extract from the SPARQL query given by the benchmarks.
Relation recall measures the proportion of golden relations edges that are successfully predicted, which reflects the method’s sensitivity to retrieve the question-relevant relations in the subgraphs:
$R_{rel}=\frac{count(\{R|R\in G_{k}\}\cap\{R|R\in G_{gold}\})}{count(\{R|R\in G_{gold}\})}.$ (10)
Minimum edit distance $D(G_{1},G_{2})$ is defined as the total number of operations required to transform one graph $G_{1}$ into another graph $G_{2}$ by sequentially adjusting its edges one by one:
[…]
When measuring $D(G_{k},G_{gold})$, the lower the distance, the smaller the structural difference between the predicted graph structure and the correct reasoning subgraph is. We combine the WebQSP and CWQ datasets and classify the test set based on the structure of the given golden subgraph with two features: number of leaf nodes and the total number of relations. The detailed statistic result can be found in Appendix E. We compare LDR with finetuned generative method RoG to evaluate the method’s performance to retrieve different types of subgraphs, the results are shown in Table 2.
Across all types of subgraph structures, it is evident that our method consistently achieves higher relation recall and lower average edit distance, which demonstrates significant enhancement of the LLM’s capability to retrieve question-related subgraphs brought by our proposed discriminative searching strategy.
Subtask Strategy WebQSP CWQ
Hits@1 F1 Graph Hits Hits@1 F1 Graph Hits
- LDR 0.840 0.845 0.884 0.802 0.820 0.894
Answer Inference Generate answer 0.761 0.744 0.884 0.684 0.679 0.894
Subgraph pruning Remove entity type 0.764 0.776 0.884 0.741 0.770 0.894
Ignore constraints 0.737 0.764 0.884 0.548 0.632 0.894
Subgraph Searching Chain-like searching 0.739 0.803 0.896 0.444 0.581 0.844
Table 3: Ablation study of the strategies to remormulate KGQA process as discriminative searching and reasoning process. Those strategies each corresponds to one of the LDR subtasks abbreviate as subgraph retrieval, subgraph pruning and answer inference.
4.4 Reduction of Redundant Information
Figure 3: The average size (number of triplets) of retrieved subgraphs.
Following the analysis of subgraph quality, we calculated the average size of the retrieved subgraphs. As shown in Figure 3, there is a significant reduction in the average size of the retrieved subgraphs, indicating that the LDR method effectively improves the efficiency of subgraph retrieval by recalling fewer but higher-quality subgraph triples.
Figure 4: The trend of average Hits@1 as the size (number of triplets) of retrieved subgraph increases.
Redundant information may significantly increase the context length, thereby affecting the performance of the generative reasoning process. We analyze the impact of the size of the retrieved subgraph (i.e., the number of triples included) on the overall performance of the strategy. In addition to using RoG with finetuned chain-of-thought reasoning, we implement the in-context reasoning strategy proposed by ToG with the subgraphs retrieved by LDR. To ensure fairness, we use Llama2_7b as the base model for all experiments. As shown in Figure 4, as the number of recalled subgraph triples increases, the performance of generative reasoning methods declines, whereas the strategy adopted by LDR remains more stable. This observation suggests that generation-based reasoning strategies are more sensitive to the size of the subgraphs compared to the discriminative strategy employed by LDR.
4.5 Mitigation of Reasoning Hallucinations
Figure 5: Case frequency of different types of errors.
To analyze the effect of adopting discriminative strategies on ungrounded reasoning with hallucinatory behavior, we collected and examined the frequency of error cases in LDR. Since we decompose the KGQA process into three subtasks executed sequentially, we can categorize all errors into the following three types: 1) E1—failed subgraph searching; 2) E2—insufficient subgraph pruning; 3) E3—incorrect answer location (refer to Appendix D for more details). We analyzed the frequency of these different error types, and the results are shown in Figure 5.
According to the result, compared to the generation-based method, LDR significantly reduces the frequency of E3 errors on both benchmarks and also reduces the overall frequency of E2 errors, which proves that LDR alleviates ungrounded reasoning behaviors in most cases by applying discriminative reasoning strategies. The dropped case frequency of E1 is consistent with our claim that the LDR enhances subgraph searching capability of the LLM.
4.6 Ablation Study of LDR
To validate the effectiveness of the discriminative strategies adopted in LDR, we ablate the strategies we adopted for three subtasks in LDR and observe the changes in performance: 1) Generate answer: ablate our discriminative answer inference strategy and let the model generate the answer based on subgraph $G_{k}$; 2) Remove entity type: replace semantic nodes in $S_{k}$ with unified names "entity_id"; 3) Ignore constraints: skip the subgraph pruning process; 4) Chain-like searching: forbid branching in the subgraph searching subtask, which is similar to the strategy used in RoG. We use "Graph Hits" to measure the proportion of retrieved subgraphs containing at least one correct answer:
$Graph\_Hits=\begin{cases}1,ifA\cap G_{k}\neq\emptyset\\ 0,ifA\cap G_{k}=\emptyset\end{cases}$ (12)
The results are shown in Table 3, which validates the effectiveness of all the LDR strategies.
Base Model WebQSP CWQ
Hits@1 F1 Hits@1 F1
Vicuna-7b 0.809 0.828 0.778 0.794
Llama-7b 0.830 0.842 0.799 0.823
Llama2-7b 0.840 0.845 0.802 0.820
Llama3-8b 0.827 0.845 0.812 0.831
Qwen2.5-7b 0.825 0.840 0.809 0.821
Table 4: We change the base model and report the performances. Replacing base models does not have significant impact on the performance of LDR.
0010101010252525255050505075757575100100100100000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8proportion of training data (%)ReasonSearch
(a) WebQSP
00101010102525252550505050757575751001001001000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8proportion of training data (%)ReasonSearch
(b) CWQ
Figure 6: The Hits@1 performance using different proportion of finetuning data.
4.7 Further Analysis
Model Universality of LDR. To analyze the model Universality of LDR, we test the performance of the LDR method based on different backbone models, and the results are shown in Table 4. The results indicate that changing the base model has no significant impact on the method’s performance, highlighting the universality of the LDR approach.
Utilization Efficiency of Training Data. We examine the utilization efficiency of training data in each LDR subtask, with the results shown in Figure 6. We combine subgraph pruning and answer inference into a single reasoning component and fine-tune two separate models from scratch using Llama2-7b. When evaluating the performance of one model, we use the other model in its fully fine-tuned form. Both models require only about 50% of the training data to achieve performances close to those models trained with entire dataset. Moreover, in terms of data requirements among subgraph retrieval, subgraph pruning and answer inference, subgraph retrieval demands more data and poses greater challenges to the LLM.
Case Study. We present cases of solving KGQA problems using the LDR method in Appendix G. LDR provides effective explicit intermediate reasoning information, which adds to the explainability of the overall KGQA process.
5 Related Works
One of the challenges in the research of KGQA task is to determine how to enable models to learn reasoning processes over knowledge graphs ( Zhu et al., 2024). Early approaches attempted to use networks such as key-value memory networks or graph neural networks to learn representations of inference paths Miller et al. (2016); Yasunaga et al. (2021)). Other works tried to teach models to construct database queries for knowledge graphs (such as SPARQL), allowing them to directly retrieve answers from the graph Gu and Su (2022); Ye et al. (2022)). However, with the advent of large language models (LLMs), the focus has shifted towards leveraging the reasoning capabilities of these models for knowledge inference. For example, UniKGQA ( Jiang et al., 2022) unifies the graph retrieval and reasoning process into a single model using LLMs.
Recently, a popular trend has been the design of effective interactive tools that enable models to engage with knowledge graphs to complete the reasoning process step-by-step, as seen in approaches like KGAgent (Jiang et al., 2024). LLMs have demonstrated surprising capabilities in long-horizon planning and reasoning (Zhong et al., 2024; Wang et al., 2024). By designing a structured interaction process, ToG has achieved remarkable zero-shot reasoning abilities on several KGQA benchmarks with the assistance of GPT-4 (Sun et al., 2024a). FAME (Hong et al., 2023) introduces Monte Carlo planning to generate reliable reasoning steps, while KD-CoT (Wang et al., 2023a) retrieves relevant knowledge from KGs to produce faithful reasoning plans for LLMs. RoG employs chain-like subgraph planning and distills GPT-4’s CoT (Chain-of-Thought) reasoning capability to achieve reliable reasoning processes over knowledge graphs, achieving state-of-the-art performance (LUO et al., 2024).
6 Conclusion
In this paper, we propose a novel LLM-based discriminative reasoning method, LDR, to address the existing issues of generation-based KGQA methods. Experimental results show that by adopting discriminative strategies, the LDR method not only enhances the capability of LLMs to retrieve question-related subgraphs but also alleviates ungrounded reasoning caused by hallucinations in the generation process.
Limitations
In the future, we will focus on the following directions to further extend the current work: 1) Conduct analysis and research on KGQA reasoning behaviors and existing issues using scaled models, including GPT-4; 2) Although this work has improved interaction efficiency compared to previously proposed interactive methods (see Appendix H), exploring efficient interaction methods between large models and knowledge graphs remains a valuable research direction; 3) Research on explainable KGQA reasoning processes at the step level: Based on a clear reasoning process, future work can further develop KGQA methods with fine-grained explanations at the level of operational steps, enhancing the explainability and reliability of the KGQA process.
Appendix A Semantic Entity Types
Here we demonstrate different semantic entity types in Table 5.
Type Definition Example
entity
Real entities include
[…]
and so on
Micheal
topic
Topic id entities which
is used to connect
entities with the
same topic, its id
has no actual meanings
m.01428y
num Numbers 240.15
date Dates 2015\08\10
Table 5: Entity types with its definition and example
Appendix B Dataset preprocessing details
Due to the huge volume of established Freebase knowledge graph, directly interacting with Freebase through SPARQL is inefficient and may result in unnecessary syntax errors. Following UniKGQA (Jiang et al., 2022), we extract subgraphs from Freebase using breadth-first search for each question, which are then used for the subgraph searching process. Additionally, we expand these subgraphs using the SPARQL queries provided in the benchmarks to ensure the presence of constraint branches.
SPARQL queries contain the subgraph information necessary to complete a comprehensive graph query. These queries are composed of graph structure triples and filtering conditions. In WebQSP, CWQ, and most KGQA datasets built on Freebase, each question corresponds to a specific SPARQL query. Therefore, the paths included in SPARQL effectively represent the correct subgraph structure required to answer the current question.
In previous work, subgraphs obtained using shortest path search methods typically formed chain-like structures. Compared to the information contained in SPARQL, these structures: 1) might not be logically coherent search paths, and 2) could miss some branches on certain nodes along the path. To enable our method to proceed smoothly, we extracted additional subgraph structures with all possible branches related to the question from Freebase based on the structural information inherent in the SPARQL queries. These were added to the original dataset (for a reference to the original dataset, see RoG). The specific implementation logic can be found in the corresponding functions in the open-source code, and will not be elaborated here.
Appendix C Metrics
Here we outline the metrics calculation formulas and their corresponding meanings that were not detailed in the main text.
Hits@1 Hits@1 calculates the proportion of questions for which the first answer given by the model is correct. Given $A_{pre}$ is the predicted list of answers, and $A_{gold}$ is the list of golden answers, $A_{pre}[0]$ as the very first answer the model predict, then we have:
$Hits@1=\frac{count(A_{pre}[0]\in A_{gold})}{count(questions)}.$ (13)
For example, if the correct answer is "apple" and the model answers "pear, apple, banana," then Hits@1 for this question is 0. It is important to note that this metric can sometimes be miscalculated as follows:
$Hits@1=\frac{count(A_{pre}\cap A_{gold}\neq\emptyset)}{count(questions)}.$ (14)
With this incorrect calculation, the Hits@1 would be higher. For the above example, the Hits@1 for this question would incorrectly become 1.
Recall, precision and F1 We adopt the same calculation method as previous work, using the Macro-F1 scoring method. First, we calculate the precision and recall for each test sample. Then, we average them based on the number of samples to obtain the overall recall and precision. Finally, we use the harmonic mean of the overall recall and precision to calculate the overall F1 score.
Appendix D Types of Errors
Does retrieved
subgraph contains correct answer?
Is the very first answer predicted correct? Case type
Yes Yes C1
No, but the correct answer exist in the predicted list E2
No, and there is no correct answer in the predicted list E3
No Yes C2
No E1
Table 6: Error Case type definitions.
Case Type CWQ WebQSP
RoG Total Seperate Total Seperate
C1 1645 1208
E2 2390 132 1363 99
E3 613 56
[…]
E1 1057 693 257 179
LDR
C1 2784 1342
E2 154 73
E3 3049 111 1449 34
[…]
E1 398 398 171 171
Table 7: Frequency count of different cases.
Leaf Count Edge count in the subgraph Total Percentage
1 2 3 4 5 6 7 8
2 921 1453 1267 400 22 0 0 0 4063 0.802
3 0 0 278 217 208 5 0 0 708 0.139
4 0 0 0 71 41 44 2 3 161 0.031
5 0 0 0 0 34 40 57 0 131 0.025
Table 8: Statistics of questions with different knowledge subgraph structure. This is the statistic result combining WebQSP and CWQ’s test set.
Tree Search Stage Prompt Template
Below is a wikipedia question,
you can retrieve a graph to help you answer the question.
The retrieved graph information is given as information triple like (Entity1, Relation, Entity2),
or only the name of the start entity.
Decide which entity and corresponding relation to retrieve next,
response in form of ’entity+relation’.
Response ’None’ if the retrieved graph is
informative enough to answer the question.
Question:
<The Question>
Retrieved graph: <The retrieved subgraph structure $S_{k}$, in forms of triples>
Next retrieve:
Tree Pruning Stage Prompt Template
Locate Constraints:
Below is a question with a support graph presented as
triples (entity A, relation, entity B).
The entity name in the support graph is ’type_id’.
’Type’ denotes the entity type, which includes four types:
ordinary entity (entity), topic entity (topic),
number (num), and date (date).
’Id’ is an incremental identifier used to distinguish entities.
Please match all the constrains with one of the entity in the support graph.
Support graph:
<The retrieved subgraph structure $S_{k}$, in forms of triples>
Question:
<The Question>
Constraints:
<List of constraint entities>
Determine result:
Locate Answer:
Below is a question with a support graph presented as
triples (entity A, relation, entity B).
The entity name in the support graph is ’type_id’.
’Type’ denotes the entity type, which includes four types:
ordinary entity (entity), topic entity (topic),
number (num), and date (date).
’Id’ is an incremental identifier used to distinguish entities.
Please select the answer from the support graph by choosing the right entity.
Support graph:
<The retrieved subgraph structure $S_{k}$, in forms of triples>
Question:
<The Question>
Answer entity:
Table 9: Prompt Template use in LDR
We categorize the answering process into five scenarios, with three of these ultimately resulting in incorrect answers. We present the overall definitions in Table 6 and the frequency statistics for the five scenarios in Table 6. Here we further explain the definition of the three error cases.
E1 (failed subgraph searching) is directly related to the graph search ability of the model. If the answer is not included in the retrieved subgraph, the model can not actually obtain the answer through inference and pruning.
We attribute E2 to lack of subgraph pruning as it indicates the presence of incorrect answer entities at the selected answer position. We detect E2 as the cases when the first answer is wrong all the correct answers are listed after the The lack of pruning may be caused by: 1) Omission of branches in the structure, which means the LLM fails to retrieve necessary entities; 2) Failure on matching constraints with the correct position. To avoid such errors, the model should have stronger searching and constraint locating capabilities.
We attribute E3 to wrong answer location since the answer list contains no golden answer. Although generation-based methods generates the answer rather than selecting the position, the inability to infer the answer from the graph containing the correct answer is considered as a similar location error. To avoid such errors, the model should have stronger subgraph reasoning and answer positioning capabilities.
The previous generation-based method can sometimes provide the correct answer even when the subgraph does not contain the correct answer, whereas LDR does not exhibit this behavior (see case C2 in Table 7).
Appendix E Statistics of Subgraph Structure
After categorizing questions based on the structure of their corresponding knowledge subgraphs, we count the number of questions in each class(see Figure 8), and find that there is a relative scarcity of graph-structured data with single or multiple branches.
Many questions with leaf count 2 is free from constraints, while these issues make up the vast majority(80.2%) of the test set. This proportional relationship also appears in the training set, which means the model will see more simple graph structures during training process. This may lead the model to prematurely halt the search by favoring structures with fewer branches. However, introducing minimum branching threshold to force the LLM to search more branches before it terminates the search stage may obstacle normal tree search behavior (see Table 11). This remains a topic worth to be discussed in the future.
Case 1
Question:
what does jamaican people speak?
Tree search stage output $S_{k}$:
(’Jamaica’, ’location.country.languages_spoken’, ’entity_1’)
Real subgraph $G_{k}$ retrieve from Freebase:
(’Jamaica’, ’location.country.languages_spoken’, ’Jamaican English’)
(’Jamaica’, ’location.country.languages_spoken’, ’Jamaican Creole English Language’)
Tree pruning stage output:
No constrain, Answer is “entity_1”
LDR Output:
’Jamaican English’, ’Jamaican Creole English Language’
Case 2
Question:
who plays the voice of kitt in knight rider?
Tree search stage output $S_{k}$:
("Knight Rider", "tv.tv_program.regular_cast", "topic_1")
("topic_1", "tv.regular_tv_appearance.character", "entity_2")
("topic_1", "tv.regular_tv_appearance.actor", "entity_3")
Real subgraph $G_{k}$ retrieve from Freebase:
("Knight Rider", "tv.tv_program.regular_cast", "m.03lj4m5")
("m.03lj4m5", "tv.regular_tv_appearance.character", "KITT")
("m.03lj4m5", "tv.regular_tv_appearance.actor", "William Daniels")
("m.02h9cb0", "tv.regular_tv_appearance.actor", "David Hasselhoff")
("m.03lj4m0", "tv.regular_tv_appearance.actor", "Edward Mulhare")
("m.03lj4lk", "tv.regular_tv_appearance.actor", "Patricia McPherson")
Tree pruning stage output:
Constraint: "entity_2= KITT", Answer is “entity_3”
LDR Output:
’William Daniels’
Case 3
Question:
where did kevin love go to college?
Tree search stage output $S_{k}$:
("Kevin Love", "people.person.education", "topic_1")
("topic_1", "education.education.institution", "entity_2")
("entity_2", "common.topic.notable_types", "entity_3")
Real subgraph $G_{k}$ retrieve from Freebase:
("Kevin Love", "people.person.education", "m.04ftwdc")
("Kevin Love", "people.person.education", "m.045z4gx")
("m.04ftwdc", "education.education.institution", "University of California, Los Angeles")
("m.045z4gx", "education.education.institution", "Lake Oswego High School")
("University of California, Los Angeles", "common.topic.notable_types", "College/University")
("Lake Oswego High School", "common.topic.notable_types", "School")
Tree pruning stage output:
Constraint: "entity_3 = College/University", Answer is “entity_2”
LDR Output:
’University of California, Los Angeles’
Table 10: Case Study of LDR
Leaf Node Number 2 3 4 5
with threshold 0.847 0.741 0.625 0.604
w/o threshold 0.845 0.750 0.636 0.738
Table 11: The average Hits@1 performance on questions with different subgraph structures. Manually add minimum branch threshold during tree search process. The performance drops as we manually add the threshold.
Appendix F Prompt Templates
We demonstrate all the prompt templates used in LDR in Table 9, including the template for tree searching, locating constraints and locating answer entity.
Appendix G Case Study
We present two clear process examples of conducting KGQA tasks using LDR in Table 10.
Appendix H Reduction Interaction Overhead
The process of interactive analysis between LLMs and knowledge graphs is quite time-consuming, especially when the extremely long context brought by large subgraphs further delays reasoning efficiency. However, by improving search accuracy, this paper has significantly reduced the number of interactions compared to ToG Sun et al. (2024a). As shown in Table 12, LDR adopts an efficient greedy search strategy, which has halved the overall number of model calls.
WebQSP CWQ
ToG 11.2 14.3
LDR 3.9 5.7
Table 12: The average number of LLM calls.